OpenStudy (anonymous):

Solve : 5e^-x -7=2 Please Show Work !

OpenStudy (anonymous):

\[5e^{-x}-7=2\] add \(7\) get \[5e^{-x}=9\] divide by \(5\) and get \[e^{-x}=\frac{9}{5}\]so the first two steps are familiar elementary algebra in "solving for \(x\)"

OpenStudy (anonymous):

now comes the only non algebra step, rewrite in logarithmic form \[e^{-x}=\frac{9}{5}\iff -x=\ln(\frac{9}{5})\]

OpenStudy (anonymous):

and finally back to algebra \[-x=\ln(\frac{9}{5})\iff x=-\ln(\frac{9}{5})\]

OpenStudy (anonymous):

if you want to be fancy, you can write \[-\ln(\frac{9}{5})=\ln(\frac{5}{9})\] or even \[\ln(5)-\ln(9)\]

OpenStudy (anonymous):

5e^-x=9 e^-x=9/5 x=In(5/9)